OpenStudy (anonymous):

In the below system, solve for y in the first equation. x - 3y = -6 2x - 7y = 10 1/3x + 2 -1/3x + 6 -x + 2 -1/3x + 2

OpenStudy (phi):

x - 3y = -6 if you add -x to both sides -x + x - 3y = -6 -x can you simplify -x + x on the left side ?

OpenStudy (anonymous):

just x right?

OpenStudy (phi):

-x + x means the same as x - x you have 1 x and take away an x. How many x's ?

OpenStudy (phi):

x - 3y = -6 if you add -x to both sides -x + x - 3y = -6 -x the -x + x is 0 x's (no x's), so we can simplify this to -3y = -x -6 ( put the -x first, because the answers have x first) now divide both sides by -3 that means EVERY TERM on both sides write /-3 under each term what do you get ?